### v2

# Symplectic fillings of unit cotangent bundles of spheres and applications

Abstract.
We prove the uniqueness, up to diffeomorphism, of symplectically aspherical fillings of the unit cotangent bundle of odd-dimensional spheres. As applications, we first show the non-existence of exact symplectic cobordisms between some $5$-dimensional Brieskorn manifolds. We also determine the diffeomorphism types of closed symplectic $6$-manifolds with certain codimension $2$ symplectic submanifolds.
1. Introduction
This paper is concerned with the topology of symplectic fillings of a given contact manifold. Based on our interest, throughout the paper, symplectic filling always means strong symplectic filling. We recall that a strong symplectic filling of a (cooriented) contact manifold $(Y,\xi)$ is a compact connected symplectic manifold $(W,\omega)$ satisfying that $\partial W=Y$, and there exists a Liouville vector field $X$ for $\omega$, defined near and pointing outward along $\partial W$, which satisfies $\xi=\operatorname{Ker}(i_{X}\omega|_{T\partial W})$ (as cooriented contact structure). We require $(W,\omega)$ to be symplectically aspherical, i.e. $\omega]|_{\pi_{2}(W)}=0$, which suffices to ensure minimality of symplectic fillings $(W,\omega)$.
Gromov [Gro85], Eliashberg [Eli90] and McDuff [McD90] independently showed that the standard contact $3$-sphere $(S^{3},\xi_{\mathrm{st}})$ has a unique symplectically aspherical filling up to diffeomorphism. The uniqueness result for $(S^{2n-1},\xi_{\mathrm{st}})$, $n\geq 3$, was proven by Eliashberg–Floer–McDuff [McD91]. The standard contact sphere $(S^{2n-1},\xi_{\mathrm{st}})$ serves as the simplest example of a subcritically Stein fillable contact manifold. Barth–Geiges–Zehmisch [BGZ19] showed the uniqueness of diffeomorphism types of symplectically aspherical fillings of a simply connected subcritically Stein fillable contact manifold $(Y,\xi)$ of $\dim Y\geq 5$.
A primary example of a critically Stein fillable contact manifold is the unit cotangent bundle $ST^{*}B$ of a closed Riemannian manifold $B$ equipped with the canonical contact structure $\xi_{\operatorname{can}}$. This contact manifold has a natural symplectic filling, namely the disk cotangent bundle $DT^{*}B$ with the canonical symplectic structure $\omega_{\operatorname{can}}$. The question is whether there exists another symplectic filling. In the case $\dim B=2$, the answer to this question is somewhat well understood. McDuff [McD90] showed that all of symplectically aspherical fillings of $(ST^{*}S^{2},\xi_{\operatorname{can}})$ are diffeomorphic to $DT^{*}S^{2}$, later which was strengthened to uniqueness up to Stein homotopy by Hind [Hin03]. For $(ST^{*}T^{2},\xi_{\operatorname{can}})$, Stipsicz [Sti02] showed that any Stein filling is homeomorphic to $DT^{*}T^{2}$, and then Wendl showed that $(DT^{*}T^{2},\omega_{\operatorname{can}})$ is the unique symplectically aspherical filling up to symplectic deformation. Even in the case where $B$ is an orientable higher-genus or non-orientable surface, topological restrictions of symplectic fillings of $(ST^{*}B,\xi_{\operatorname{can}})$ are studied: [LMY17], [SVHM17], [LO18] and [GZ24]. In contrast, for the case $\dim B\geq 3$, the topology of symplectic fillings of $(ST^{*}B,\xi_{\operatorname{can}})$ is less explored. For $B=T^{n}$, Bowden–Gironella–Moreno [BGM22] and independently Geiges–Kwon–Zehmisch [GKZ23] showed the uniqueness of symplectically aspherical fillings up to diffeomorphism. In addition, Kwon–Zehmisch [KZ19] provided a uniqueness result for the diffeomorphism types of symplectically aspherical fillings of $(ST^{*}S^{2d+1},\xi_{\operatorname{can}})$ ($d\geq 1$) , assuming the existence of a certain complex hypersurface.
The main theorem in this paper is a uniqueness result for symplectically aspherical fillings of $(ST^{*}S^{2d+1},\xi_{\operatorname{can}})$, which has also recently been obtained by Bimmermann–Stratmann–Zehmisch [BSZ25].
Theorem 1.1.
For any integer $d\geq 1$, let $(W,\omega)$ be a symplectically aspherical filling of the unit cotangent bundle $(ST^{*}S^{2d+1},\xi_{\operatorname{can}})$. Then, $W$ is diffeomorphic to $DT^{*}S^{2d+1}$.
This theorem will be proved in Section 3 after we discuss moduli spaces of holomorphic spheres in Section 2. The present proof corrects and simplifies the earlier argument in the previous version and is inspired by the techniques of [KZ19, BSZ25]. Following the approach of [KZ19, BSZ25], we construct a symplectic cap of $(ST^{*}S^{2d+1},\xi_{\operatorname{can}})$ using the Audin–Lalonde–Polterovich [ALP94] Lagrangian embedding of $S^{2d+1}$ to close off the boundary of a symplectically aspherical filling $(W,\omega)$ of $(ST^{*}S^{2d+1},\xi_{\operatorname{can}})$. We then study moduli spaces of holomorphic spheres in the resulting capped manifold. Applying the so-called degree method (see e.g. [McD91, BGZ19]) to these moduli spaces yields topological information about $(W,\omega)$, which determines its fundamental group and homology. Finally, combining these computations with a result of [KZ19] establishes Theorem 1.1. A notable feature of our argument is that the properness of an evaluation map from a moduli space can be shown more directly than in [BSZ25]; see Lemma 2.2 and Remark 2.3.
We would like to note that, in a forthcoming paper [KO], the authors prove that symplectically aspherical fillings of $(ST^{*}S^{n},\xi_{\operatorname{can}})$ are simply connected for all $n\geq 3$ by using a different method.
In Section 4, we will give two applications of the main theorem. The first application is to exact symplectic cobordisms. For each integer $\ell\geq 1$, define
$Y_{\ell}=\{(z_{1},z_{2},z_{3},z_{4})\in\mathbb{C}^{4}\mid z_{1}^{2\ell}+z_{2}^{2}+z_{3}^{2}+z_{4}^{2}=0\}\cap S^{7},$
which is the so-called Brieskorn manifold of the singularity $z_{1}^{2\ell}+z_{2}^{2}+z_{3}^{2}+z_{4}^{2}$. We equip it with the contact structure $\xi_{\ell}$ induced by the standard one on $S^{7}$. Note that $(Y_{1},\xi_{1})$ is contactomorphic to $(ST^{*}S^{3},\xi_{\operatorname{can}})$. It is known that all of the smooth manifolds $Y_{\ell}$ are diffeomorphic to $ST^{*}S^{3}$. In contrast, all of the contact manifolds $(Y_{\ell},\xi_{\ell})$ are mutually non-contactomorphic by a result of Uebele [Ueb16]. Therefore, there is no obstruction to the existence of smooth cobordisms between manifolds $Y_{\ell}$, and, moreover, there is an exact symplectic cobordism from $(Y_{\ell},\xi_{\ell})$ to $(Y_{\ell^{\prime}},\xi_{\ell^{\prime}})$ if $\ell\leq\ell^{\prime}$. However, unlike smooth cobordisms, the latter fact does not automatically guarantee the existence of a reverse exact symplectic cobordism; in fact, Kwon–van Koert [KvK16] posed a question about this existence. We partially answer this question in Theorem 4.1: there is no exact symplectic cobordism from $(Y_{\ell},\xi_{\ell})$ to $(Y_{1},\xi_{1})$ for any $\ell\geq 2$. A similar argument proves the non-existence of exact symplectic cobordisms from the exotic contact $5$-spheres studied in [Ust99] to the standard $5$-sphere $(S^{5},\xi_{\mathrm{st}})$; see Remark 4.5.
We will present another application of Theorem 1.1 in Section 4.2. McDuff [McD90] determined the symplectomorphism types of closed and connected symplectic $4$-manifolds containing symplectic spheres with non-negative self-intersection. This reveals that the existence of a certain type of symplectic submanifolds constrains the topology of ambient symplectic $4$-manifolds. We capture a similar phenomenon in dimension $6$. A codimension $2$ symplectic submanifold $\Sigma$ of a closed symplectic manifold $(M,\Omega)$ with $\Omega]\in H^{2}(M;\mathbb{Z})$ is called a symplectic hyperplane section on $(M,\Omega)$ if the cohomology class $\Omega]$ is Poincaré dual to $\Sigma]$. As the main result of Section 4.2, we prove that if a closed symplectic $6$-manifold $(M,\Omega)$ with $\Omega]\in H^{2}(M;\mathbb{Z})$ contains $(\mathbb{CP}^{1}\times\mathbb{CP}^{1},\omega_{\operatorname{FS}}\oplus\omega_{\operatorname{FS}})$ as a symplectic hyperplane section, then $M$ is diffeomorphic to a complex $3$-dimensional quadric (see Theorem 4.6). Note that replacing $(\mathbb{CP}^{1}\times\mathbb{CP}^{1},\omega_{\operatorname{FS}}\oplus\omega_{\operatorname{FS}})$ by $(\mathbb{CP}^{2},\omega_{\operatorname{FS}})$ yields a similar result for this symplectic $4$-manifold; see Remark 4.9. We would like to mention that Bǎdescu [Bad82] showed an analogous result for complex projective $3$-folds.
2. Moduli spaces of holomorphic spheres
2.1. Closed symplectic manifold containing a symplectic filling
Let $(W,\omega)$ be a symplectically aspherical filling of $(ST^{*}S^{2d+1},\xi_{\operatorname{can}})$ with $d\geq 1$. In this paper, the holomorphic spheres under consideration lie in a closed symplectic manifold which contains $(W,\omega)$ as a symplectic submanifold. Such a manifold can be constructed using Lagrangian embeddings of odd-dimensional spheres due to Audin–Lalonde–Polterovich [ALP94].
Consider the hyperplane bundle $O_{\mathbb{CP}^{d}}(1)\to\mathbb{CP}^{d}$. Its first Chern class is represented by $\omega_{\operatorname{FS}}$, where $\omega_{\operatorname{FS}}$ denotes the normalized Fubini–Study form on $\mathbb{CP}^{d}$, defined so that
$\int_{\mathbb{CP}^{1}}\omega_{\operatorname{FS}}=1.$
There exists a disk bundle $N(\mathbb{CP}^{d})\subset O_{\mathbb{CP}^{d}}(1)$ equipped with a symplectic form $\Omega_{0}$ such that it restricts to $\omega_{\operatorname{FS}}$ on the zero-section and admits a Liouville vector field, defined away from the zero-section, which is tangent to each fiber and points inward along the boundary; in particular, the boundary $\partial N(\mathbb{CP}^{d})$ is concave. Reversing the orientation of the fibers, $\partial N(\mathbb{CP}^{d})$ becomes the Boothby–Wang bundle associated with the symplectic manifold $(\mathbb{CP}^{d},\omega_{\operatorname{FS}})$; in fact, it is contactomorphic to the standard sphere $(S^{2d+1},\xi_{\operatorname{st}})$.
Consider the closed disk $D^{2d+2}(1/\sqrt{\pi})\subset\mathbb{C}^{d+1}$ of radius $1/\sqrt{\pi}$, equipped with the standard symplectic form $\omega_{\operatorname{st}}$. Using Liouville flows near $\partial D^{2d+2}(1/\sqrt{\pi})$ and $\partial N(\mathbb{CP}^{d})$, we can glue $D^{2d+2}(1/\sqrt{\pi})$ and $N(\mathbb{CP}^{d})$ symplectically to obtain the closed symplectic manifold
$(M_{0}=D^{2d+2}(1/\sqrt{\pi})\cup N(\mathbb{CP}^{d}),\,\omega_{0}).$
(See [KO23, Section 2.3] for further details on the construction of such glued symplectic manifolds.)
Now as in [KZ19] and [BSZ25], consider the Lagrangian embedding of the unit sphere $S^{2d+1}=\{z\in\mathbb{C}^{d+1}\mid\|z\|=1\}$ into $(D^{2d+2}(1/\sqrt{\pi})\times\mathbb{CP}^{d},\omega_{\operatorname{st}}\oplus\omega_{\operatorname{FS}})$, introduced by [ALP94], which is defined by
(2.1) $\varphi\colon S^{2d+1}\to D^{2d+2}(1/\sqrt{\pi})\times\mathbb{CP}^{d},\quad\varphi(z)=(z/\sqrt{\pi},(\overline{z})),$
where $\omega_{\operatorname{st}}\oplus\omega_{\operatorname{FS}}$ denotes the symplectic form $\mathrm{pr}_{1}^{*}\omega_{\operatorname{st}}+\mathrm{pr}_{2}^{*}\omega_{\operatorname{FS}}$ on $D^{2d+2}(1/\sqrt{\pi})\times\mathbb{CP}^{d}$, where $\mathrm{pr}_{i}$ is the projection to the $i$-th factor. Since the inclusion $\iota\colon D^{2d+2}(1/\sqrt{\pi})\times\mathbb{CP}^{d}\hookrightarrow M_{0}\times\mathbb{CP}^{d}$ satisfies $\iota^{*}(\omega_{0}\oplus\omega_{\operatorname{FS}})=\omega_{\operatorname{st}}\oplus\omega_{\operatorname{FS}}$, the image $L\coloneqq\varphi(S^{2d+1})$ is still Lagrangian in $(M_{0}\times\mathbb{CP}^{d},\omega_{0}\oplus\omega_{\operatorname{FS}})$. Take a Weinstein tubular neighborhood $N(L)$ of the Lagrangian sphere $\varphi(S^{2d+1})$ in $M_{0}\times\mathbb{CP}^{d}$ and replace it by $(W,\omega)$ symplectically. We denote the resulting glued symplectic manifold by
$(X=((M_{0}\times\mathbb{CP}^{d})\setminus N(L))\cup W,\Omega).$
Letting $\Sigma$ be the image of the zero-section of $\mathcal{O}_{\mathbb{CP}^{d}}(1)\to\mathbb{CP}^{d}$, we may assume that $X$ contains $\Sigma\times\mathbb{CP}^{d}\subset M_{0}\times\mathbb{CP}^{d}$ by choosing $N(\varphi(S^{2d+1}))$ sufficiently thin.
2.2. The moduli spaces
To define moduli spaces of holomorphic spheres, we take an $\Omega$-tame almost complex structure $J$ on $X$ which agrees with the integrable almost complex structure $J_{0}$ on $N(\mathbb{CP}^{d})\times\mathbb{CP}^{d}$ inherited from $\mathcal{O}_{\mathbb{CP}^{d}}(1)\times\mathbb{CP}^{d}$. Take a point $p_{0}\in\Sigma$. Write $C$ for the second homology class of $X$ represented by a generic complex line in $\Sigma$.
Now consider the moduli space $\widetilde{\mathcal{M}}$ consisting of all (parametrized) $J$-holomorphic spheres $u\colon\mathbb{CP}^{1}\rightarrow X$ such that $u]\coloneqq u_{*}([\mathbb{CP}^{1}])=C\in H_{2}(X;\mathbb{Z})$ and $u(0)\in\{p_{0}\}\times\mathbb{CP}^{d}$:
$\widetilde{\mathcal{M}}=\left\{\begin{gathered}u\in W^{1,p}(\mathbb{CP}^{1},X)\mid u\textrm{ is $J$-holomorphic, }[u]=C\textrm{ and }u(0)\in\{p_{0}\}\times\mathbb{CP}^{d}\end{gathered}\right\}.$
The group $G=\mathrm{Aut}(\mathbb{CP}^{1};0)$ of automorphisms of $\mathbb{CP}^{1}$ fixing 0 freely acts on $\widetilde{\mathcal{M}}$ by
$\phi\cdot u\coloneqq u\circ\phi^{-1}$
for $\phi\in G$ and $u\in\widetilde{\mathcal{M}}$; similarly, $G$ also acts on $\widetilde{\mathcal{M}}\times\mathbb{CP}^{1}$ by
$\phi\cdot(u,z)\coloneqq(u\circ\phi^{-1},\phi(z)).$
Note that for $(u,z)\in\mathcal{M}\times\mathbb{CP}^{1}$, the marked point $z$ may agree with 0. Set
$\mathcal{M}=\widetilde{\mathcal{M}}/G\quad\mathrm{and}\quad\mathcal{U}=(\widetilde{\mathcal{M}}\times\mathbb{CP}^{1})/G.$
We write $u$ for its equivalence class in $\mathcal{M}$ whenever no ambiguity arises.
Lemma 2.1.
The moduli spaces $\mathcal{M}$ and $\mathcal{U}$ are smooth manifolds of dimension $4d$ and $4d+2$, respectively, for a generic $J$.
Proof.
Since the $G$-actions on $\widetilde{\mathcal{M}}$ and $\widetilde{\mathcal{M}}\times\mathbb{CP}^{1}$ are free and proper (see e.g. [Kes10, Lemma 3.1]), the lemma follows immediately from the quotient manifold theorem once we verify that $\widetilde{\mathcal{M}}$ is a smooth manifold of dimension $4d+4$. Note that $\dim G=4$.
Let us first compute the expected dimension of $\widetilde{\mathcal{M}}$. Over the image of $\mathbb{CP}^{1}=\mathbb{CP}^{1}\times\{\mathrm{pt}\}$ under the inclusion $\mathbb{CP}^{1}\times\{\mathrm{pt}\}\hookrightarrow\Sigma\times\mathbb{CP}^{d}\hookrightarrow X$, the tangent bundle $TX$ splits into
$TX|_{\mathbb{CP}^{1}}\cong\$$(N_{(\Sigma\times\mathbb{CP}^{d})/X}\oplus T(\Sigma\times\mathbb{CP}^{d}))|_{\mathbb{CP}^{1}}\cong\mathcal{O}_{\mathbb{CP}^{1}}(1)\oplus T\Sigma|_{\mathbb{CP}^{1}}$
[…]
We obtain the same splitting over the sphere $\{\mathrm{pt}\}\times\mathbb{CP}^{1}$. Thus,
$c_{1}(TM)([u])=2(dc_{1}(\mathcal{O}_{\mathbb{CP}^{1}}(1))+c_{1}(\mathcal{O}_{\mathbb{CP}^{1}}(2)))(C)=d+2,$
and the expected dimension of $\widetilde{\mathcal{M}}$ is given by
$2c_{1}(TM)([u])+\frac{\dim X}{2}\cdot\chi(\mathbb{CP}^{1})-\mathrm{codim}(\{p_{0}\}\times\mathbb{CP}^{d})=4d+4.$
Here, $\chi(\mathbb{CP}^{1})$ denotes the Euler characteristic of $\mathbb{CP}^{1}$ and $\mathrm{codim}(\{p_{0}\}\times\mathbb{CP}^{d})$ denotes the codimension of the $2d$-dimensional submanifold $\{p_{0}\}\times\mathbb{CP}^{d}$ in $X$.
Since the algebraic intersection number between $C$ and $\Sigma\times\mathbb{CP}^{d}$ is $1$, positivity of intersections shows any holomorphic sphere belonging to $X$ is simply-covered. In light of [MS04, Lemma 3.3.1], the splitting (2.2) implies that every element $u$ of $\widetilde{\mathcal{M}}$ whose image is contained in a neighborhood of $\Sigma\times\mathbb{CP}^{d}$ is Fredholm regular. Moreover, [Wen10, Theorem 4.4.3] guarantees the existence of a generic extension $J$ of $J_{0}$ to make the other $J$-holomorphic spheres all Fredholm regular with respect to $J$. Therefore, for such a generic $J$, [MS04, Theorem 3.4.1] concludes that $\widetilde{\mathcal{M}}$ is a smooth manifold of dimension $4d+4$. ∎
Consider the following commutative diagram of smooth manifolds:
[…]
where the horizontal arrows are the quotient maps by the $G$-actions, $\mathrm{pr}_{1}$ denotes the first projection and
$\mathfrak{f}\colon\mathcal{U}\longrightarrow\mathcal{M}.$
is the map induced by $\mathrm{pr}_{1}$. In view of the freeness and properness of the $G$-actions, it follows from a standard argument in the theory of smooth manifolds that the map $\mathfrak{f}$ is smooth and, in fact, a $\mathbb{CP}^{1}$-bundle.
In other words, $\mathcal{U}$ serves as the universal family over $\mathcal{M}$. We now point out that $\mathfrak{f}$ admits two distinguished sections, which will play an important role in the next section. Consider the evaluation map
$\operatorname{ev}\colon\mathcal{U}\longrightarrow X,\qquad\operatorname{ev}([u,z])=u(z).$
The first section of the $\mathbb{CP}^{1}$-bundle is
$\operatorname{ev}^{-1}(\{p_{0}\}\times\mathbb{CP}^{d}).$
For the second section, take a generic holomorphic section of $\mathcal{O}_{\mathbb{CP}^{d}}(1)\to\mathbb{CP}^{d}$ whose image lies in $N(\mathbb{CP}^{d})$ and is disjoint from $p_{0}\in\Sigma$. Denote the image of this section by $\Sigma^{\prime}$. Then
$\operatorname{ev}^{-1}(\Sigma^{\prime}\times\mathbb{CP}^{d})$
defines another section. These are indeed sections, as follows from positivity of intersections together with the fact that the algebraic intersection numbers of $\Sigma\times\mathbb{CP}^{d}]=[\Sigma^{\prime}\times\mathbb{CP}^{d}]$ with $C$ are both equal to $1$. Moreover, since $p_{0}$ and $\Sigma^{\prime}$ are disjoint, the two sections constructed above are disjoint as well.
2.3. Compactness of the moduli spaces
Here we show the compactness of the moduli spaces.
Lemma 2.2.
The moduli spaces $\mathcal{M}$ and $\mathcal{U}$ are compact. In particular, the evaluation map $\operatorname{ev}\colon\mathcal{U}\to X$ is proper.
Proof.
It suffices to show the compactness of $\mathcal{M}$. The following argument is inspired by [KZ19, Section 3.4]. Let $(u_{\nu})$ be a sequence of holomorphic spheres $u_{\nu}\in\mathcal{M}$. By Gromov compactness, passing to a subsequence, $(u_{\nu_{k}})$ converges to a stable holomorphic sphere
$u_{\infty}=(u_{\infty}^{1},\ldots,u_{\infty}^{m})\in\overline{\mathcal{M}},$
where $\overline{\mathcal{M}}$ is the Gromov compactification of $\mathcal{M}$.
We first prove that exactly one irreducible component among $u_{\infty}^{j}$ ($j=1,\ldots,m$) is non-constant. The proof proceeds by contradiction, using symplectic energy. Assume that $m\geq 2$. We may suppose that all $u_{\infty}^{j}$ are non-constant. By the Mayer–Vietoris exact sequence together with the simply-connectedness of $ST^{*}S^{2d+1}=\partial W$, the map
$H_{2}(W;\mathbb{Z})\oplus H_{2}((M_{0}\times\mathbb{CP}^{d})\setminus N(L);\mathbb{Z})\to H_{2}(X;\mathbb{Z})$
is surjective. Since $(W,\omega)$ is symplectically aspherical, the elements of $H_{2}(W;\mathbb{Z})$ have zero symplectic energy and hence contribute nothing to the energy. Similarly, one finds that the map
$H_{2}(N(L);\mathbb{Z})\oplus H_{2}((M_{0}\times\mathbb{CP}^{d})\setminus N(L);\mathbb{Z})\to H_{2}(M_{0}\times\mathbb{CP}^{d};\mathbb{Z})$
is surjective. When $d>1$, this map is in fact an isomorphism since $H_{2}(ST^{*}S^{2d+1};\mathbb{Z})=0$. Thus, $H_{2}((M_{0}\times\mathbb{CP}^{d})\setminus N(L);\mathbb{Z})$ is generated by $C$ and $D$, where $D$ is represented by a complex line in $\{\mathrm{pt}\}\times\mathbb{CP}^{d}$. When $d=1$, in addition to those generators, $H_{2}((M_{0}\times\mathbb{CP}^{d})\setminus N(L);\mathbb{Z})$ has a fiber class of the $S^{2}$-bundle
$ST^{*}S^{3}\cong S^{3}\times S^{2}\to S^{3}$
as observed in [KZ19, Section 3.4.3]. However, since the symplectic area of this class is 0, it can be ignored. Therefore, we may express each homology class $u_{\infty}^{j}]$ as
$[u_{\infty}^{j}]=a_{j}C+b_{j}D+S_{j},$
where $a_{j},b_{j}\in\mathbb{Z}$ and $S_{j}$ is the image of an element of $H_{2}(W;\mathbb{Z})$.
By positivity of intersections,
$[u_{\infty}^{j}]\cdot[\Sigma\times\mathbb{CP}^{d}]=a_{j}\geq 0.$
Since $C\cdot[\Sigma\times\mathbb{CP}^{d}]=1$, it follows that eaxctly one $a_{j}$ equals $1$ while the others vanish. Without loss of generality, assume this occurs for $j=1$:
$[u_{\infty}^{1}]=C+b_{1}D+S_{1}\quad\textrm{ and }\quad[u_{\infty}^{j}]=b_{j}D+S_{j}\textrm{ for }j=2,\ldots,m.$
As $u_{\infty}]=[u_{\nu}]=C$, we obtain
$\Omega([u_{\infty}])=\Omega(C)=1.$
On the other hand, $u_{\infty}]=[u_{\infty}^{1}]+\cdots+[u_{\infty}^{m}]=C+(\sum_{j=1}^{m}b_{j})D+\sum_{j=1}^{m}S_{j}$, so
$\Omega([u_{\infty}])=\Omega\left(C+\left(\sum_{j=1}^{m}b_{j}\right)D+\sum_{j=1}^{m}S_{j}\right)=1+\sum_{j=1}^{m}b_{j}.$
Thus, $\sum_{j=1}^{m}b_{j}=0$, that is,
(2.3) $b_{1}=-\sum_{j=2}^{m}b_{j}.$
Since each $u_{j}^{\infty}$ is a non-constant holomorphic sphere, its symplectic energy is positive, which implies
$b_{1}>-1\quad\textrm{ and }\quad b_{j}>0\textrm{ for }j=2,\ldots,m.$
Combing this with the inequality (2.3) yields
$-1<b_{1}<0,$
contradicting the integrality of $b_{1}$. Hence $u_{\infty}$ has a unique non-constant component.
Finally we check that the limit $u_{\infty}=(u_{\infty}^{1},\ldots,u_{\infty}^{m})$ has no constant component, and therefore $u_{\infty}\in\mathcal{M}$. Suppose on the contrary that a constant component exists. Since $u_{\infty}$ has only one non-constant component, this forces the existence of a constant component with at least two marked points besides nodal points. This is impossible by the definition of the moduli space $\mathcal{M}$, which completes the proof. ∎
Remark 2.3.
In [KZ19] and [BSZ25], the ambient space $\mathbb{CP}^{1}\times\mathbb{C}^{d}\times\mathbb{CP}^{d}$ is obtained by the one-point compactification of the first $\mathbb{C}$-factor of $\mathbb{C}^{d+1}\times\mathbb{CP}^{d}$, while our $M_{0}\times\mathbb{CP}^{d}$ is obtained by capping off the $\mathbb{C}^{d+1}$-factor of $\mathbb{C}^{d+1}\times\mathbb{CP}^{d}$ using a compactifying divisor. These make the difference between [KZ19, BSZ25] and ours to show the properness of the evaluation map: in [KZ19, BSZ25], their proof crucially relies on the existence of two hypersurfaces with desired intersection behavior, one of which is demanding to find; on the other hand, our proof uses only one canonical hypersurface $\Sigma\times\mathbb{CP}^{d}$.
Remark 2.4.
The moduli space $\mathcal{M}$ a priori is not connected; neither is $\mathcal{U}$ as it is a $\mathbb{CP}^{1}$-bundle over $\mathcal{M}$. Still, all holomorphic spheres contained in $\Sigma\times\mathbb{CP}^{d}$ belong to the same connected component of $\mathcal{M}$. Hence, taking this connected component, we assume $\mathcal{M}$ and $\mathcal{U}$ to be connected throughout this paper.
3. Topology of symplectic fillings
3.1. Degree of the evaluation map
Let $(W,\omega)$ be a symplectically aspherical filling of $(ST^{*}S^{2d+1},\xi_{\operatorname{can}})$ with $d\geq 1$ and $(X,\Omega)$ the closed symplectic manifold constructed as in Section 2.1. We begin this section by analyzing the evaluation map $\operatorname{ev}\colon\mathcal{U}\rightarrow X$ defined to be $\operatorname{ev}([u,z])=u(z)$.
Lemma 3.1.
Given a point $q_{0}\in(\Sigma\times\mathbb{CP}^{d})\setminus(\{p_{0}\}\times\mathbb{CP}^{d})$, there exists a unique holomorphic sphere $u\in\mathcal{M}$ passing through $q_{0}$.
Proof.
Since $C\cdot[\Sigma\times\mathbb{CP}^{d}]=1$, positivity of intersections shows that the image of an element $u$ of $\mathcal{M}$ passing through $q_{0}$ must entirely lie in $\Sigma\times\mathbb{CP}^{d}$. By the choice of the almost complex structure $J$, such a holomorphic sphere is a genuine complex curve on $\Sigma\times\mathbb{CP}^{d}$.
We denote a curve in question by $u=(u_{1},u_{2})$ with complex curves $u_{1}\colon\mathbb{CP}^{1}\to\Sigma$ and $u_{2}\colon\mathbb{CP}^{1}\to\mathbb{CP}^{d}$. Since $[u]=C$, we have $u_{2}]=0$, which implies that $u_{2}$ is constant. As for $u_{1}$, this is a complex line in $\Sigma\cong\mathbb{CP}^{d}$ passing through the distinct two points. This concludes the uniqueness of holomorphic spheres in question. ∎
Proposition 3.2.
The evaluation map $\operatorname{ev}\colon\mathcal{U}\rightarrow X$ has degree $1$. It restricts to a proper degree $1$ map $\operatorname{ev}\colon\mathcal{U}\setminus\operatorname{ev}^{-1}((\Sigma^{\prime}\times\mathbb{CP}^{d}))\rightarrow X\setminus(\Sigma^{\prime}\times\mathbb{CP}^{d})$.
Proof.
Take a point $q_{0}\in\Sigma\times\mathbb{CP}^{d}$ as in the preceding lemma. By that lemma, the fiber $\operatorname{ev}^{-1}(q_{0})$ consists of a single element, namely $\operatorname{ev}^{-1}(q_{0})={[u,z]}$. In view of the splitting of $u^{*}TX$ in (2.2), it follows that $q_{0}$ is a regular value of $\operatorname{ev}$. Hence $\deg(\operatorname{ev})=1$. The assertion concerning the restricted evaluation map then follows directly. ∎
3.2. Uniqueness of symplectic fillings
Here we prove Theorem 1.1. The proof relies on the following lemma, which is essentially equivalent to a result of Kwon–Zehmisch [KZ19, Lemma 3.1].
Lemma 3.3 (Kwon–Zehmisch [KZ19, Lemma 3.1]).
If the embedding $\iota\colon\{p_{0}\}\times\mathbb{CP}^{d}\to X\setminus(\Sigma^{\prime}\times\mathbb{CP}^{d})$ is $\pi_{1}$-surjective and $H_{k}(\,\cdot\,;\mathbb{F})$-onto for all $k\in\mathbb{Z}$ and all fields $\mathbb{F}$, then $W$ is diffeomorphic to $DT^{*}S^{2d+1}$.
Proof of Theorem 1.1.
Recall that $\operatorname{ev}^{-1}(\{p_{0}\}\times\mathbb{CP}^{d})$ and $\operatorname{ev}^{-1}(\Sigma^{\prime}\times\mathbb{CP}^{d})$ give disjoint sections of $\mathfrak{f}\colon\mathcal{U}\to\mathcal{M}$. Moreover, since $\mathfrak{f}$ is a $\mathbb{CP}^{1}$-bundle, the space $\mathcal{U}\setminus\operatorname{ev}^{-1}(\Sigma^{\prime}\times\mathbb{CP}^{d})$ deformation retracts onto $\operatorname{ev}^{-1}(\{p_{0}\}\times\mathbb{CP}^{d})$. Thus we obtain a commutative diagram
${\operatorname{ev}^{-1}(\{p_{0}\}\times\mathbb{CP}^{d})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$$\scriptstyle{\hskip 14.0pt\operatorname{ev}}$${\{p_{0}\}\times\mathbb{CP}^{d}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota}$${\mathcal{U}\setminus\operatorname{ev}^{-1}(\Sigma^{\prime}\times\mathbb{CP}^{d})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{ev}}$${X\setminus(\Sigma^{\prime}\times\mathbb{CP}^{d})}$
where the vertical arrows are inclusions. As $\operatorname{ev}\colon\mathcal{U}\setminus\operatorname{ev}^{-1}(\Sigma^{\prime}\times\mathbb{CP}^{d})\to X\setminus(\Sigma^{\prime}\times\mathbb{CP}^{d})$ has degree $1$ by Proposition 3.2, it induces a surjection on $\pi_{1}$ and is onto on $H_{k}(\,\cdot\,;\mathbb{F})$ for all $k\in\mathbb{Z}$ and all fields $\mathbb{F}$. In view of the homotopy equivalence $\operatorname{ev}^{-1}(\{p_{0}\}\times\mathbb{CP}^{d})\simeq\mathcal{U}\setminus\operatorname{ev}^{-1}(\Sigma^{\prime}\times\mathbb{CP}^{d})$, the same holds for $\iota$. Therefore, Lemma 3.3 implies that $W$ is diffeomorphic to $DT^{*}S^{2d+1}$. ∎
Remark 3.4.
The above proof can be compared to the proof of the uniqueness of symplectically aspherical fillings of $(S^{2n-1},\xi_{\operatorname{st}})$ given by Ghiggini–Niederkrüger [GNE22, Corollary 5.3]. In their proof, they consider the closed symplectic manifold $(X^{\prime},\Omega^{\prime})$ obtained from a given symplectically aspherical filling $(W^{\prime},\omega^{\prime})$ of $(S^{2n-1},\xi_{\operatorname{st}})$ by capping with the disk bundle over $\mathbb{CP}^{n-1}$ associated with $\mathcal{O}_{\mathbb{CP}^{n-1}}(1)$. Then they determine the fundamental group and homology of $W^{\prime}$ using the moduli space of holomorphic spheres $\mathcal{M}^{\prime}$ in $X^{\prime}$ passing through a fixed point of $\mathbb{CP}^{n-1}\subset X^{\prime}$ and the universal family $\mathcal{U}^{\prime}$ over $\mathcal{M}^{\prime}$. Finally, they apply the h-cobordism theorem to determine the diffeomorphism type of $W^{\prime}$. We observe that the moduli spaces $\mathcal{M}^{\prime}$ and $\mathcal{U}^{\prime}$ are similar to our $\mathcal{M}$ and $\mathcal{U}$, respectively.
4. Applications
4.1. Symplectic cobordisms
Let $(Y_{+},\xi_{+}=\operatorname{Ker}(\alpha_{+}))$ and $(Y_{-},\xi_{-}=\operatorname{Ker}(\alpha_{-}))$ be (cooriented) contact manifolds. A compact exact symplectic manifold $(X,d\lambda)$ with boundary $\partial X=-Y_{-}\sqcup Y_{+}$ is called an exact symplectic cobordism from $(Y_{-},\xi_{-})$ to $(Y_{+},\xi_{+})$ if the Liouville vector field on $X$, defined to be $d\lambda$-dual to $\lambda$, points inward along $Y_{-}$ and outward along $Y_{+}$, and $\xi_{\pm}=\operatorname{Ker}(\lambda|_{TY_{\pm}})$. Here $Y_{\pm}$ and $X$ are oriented by the volume forms $\alpha_{\pm}\wedge(d\alpha_{\pm})^{n-1}$ and $(d\lambda)^{n}$, respectively, and $Y_{-}$ means $Y_{-}$ with the reversed orientation. If $(Y_{-},\xi_{-})$ is the empty set, then the exact symplectic cobordism $(X,d\lambda)$ is called an exact symplectic filling of $(Y_{+},\xi_{+})$.
Recall that for each integer $\ell\geq 1$ the $5$-manifold $Y_{\ell}$ is defined to be
$Y_{\ell}=\{(z_{1},z_{2},z_{3},z_{4})\in\mathbb{C}^{4}\mid z_{1}^{2\ell}+z_{2}^{2}+z_{3}^{2}+z_{4}^{2}=0\}\cap S^{7}.$
This is the so-called Brieskorn manifold, known to be diffeomorphic to $S^{2}\times S^{3}$ (see e.g. [DK75, Proposition 6.1]). It also carries the canonical contact structure $\xi_{\ell}$, induced by the standard contact structure on the unit $7$-sphere $S^{7}=\{z\in\mathbb{C}^{4}\mid\|z\|=1\}$. The Brieskorn contact manifold $(Y_{1},\xi_{1})$ is contactomorphic to the unit cotangent bundle $(S^{*}TS^{3},\xi_{\operatorname{can}})$ (see for example [KvK16, Lemma 3.1] and [May09, Lemma 5.3]). A Milnor fiber $V_{\ell}$ of the singularity $z_{1}^{2\ell}+z_{2}^{2}+z_{3}^{2}+z_{4}^{2}$ gives an exact symplectic filling of $(Y_{\ell},\xi_{\ell})$. According to [AGLV98, Section 2.7], for $\ell$ and $\ell^{\prime}$ with $\ell<\ell^{\prime}$, the singularity $z_{1}^{2\ell}+z_{2}^{2}+z_{3}^{2}+z_{4}^{2}$ is adjacent to $z_{1}^{2\ell^{\prime}}+z_{2}^{2}+z_{3}^{2}+z_{4}^{2}$; hence [Kea14, Lemma 9.9] leads to an exact symplectic embedding of a Milnor fiber $V_{\ell}$ into the completion of $V_{\ell^{\prime}}$. The latter is diffeomorphic to $V_{\ell^{\prime}}\cup([0,\infty)\times Y_{\ell^{\prime}})$, and the image of the embedding is contained in $V_{\ell^{\prime}}\cup([0,c)\times Y_{\ell^{\prime}})$ for some $c>0$ by compactness. Thus, truncating the completion at the hypersurface $\{c\}\times Y_{\ell^{\prime}}$ and removing the interior of the embedded $V_{\ell}$ shows the existece of an exact symplectic cobordism from $(Y_{\ell},\xi_{\ell})$ to $(Y_{\ell^{\prime}},\xi_{\ell^{\prime}})$. The following theorem tells us that this does not hold for the reverse direction in general.
Theorem 4.1.
For any $\ell>1$, there exists no exact symplectic cobordism from $(Y_{\ell},\xi_{\ell})$ to $(Y_{1},\xi_{1})$.
Proof.
Suppose for the sake of contradiction that there exists an exact symplectic cobordism $(W_{\ell,1},d\lambda)$ from $(Y_{\ell},\xi_{\ell})$ to $(Y_{1},\xi_{1})$. Then, gluing $(W_{\ell,1},d\lambda)$ with $V_{\ell}$ symplectically gives an exact symplectic filling $(W,\omega)$ of $(Y_{1},\xi_{1})$. Consider the Mayer–Vietoris exact sequence for the pair $(V_{\ell},W_{\ell,1})$:
(4.1) $H_{3}(Y_{\ell};\mathbb{Q})\rightarrow H_{3}(V_{\ell};\mathbb{Q})\oplus H_{3}(W_{\ell,1};\mathbb{Q})\rightarrow H_{3}(W;\mathbb{Q})\xrightarrow{\varphi}H_{2}(Y_{\ell};\mathbb{Q})$
We have $H_{3}(Y_{\ell};\mathbb{Q})\cong H_{2}(Y_{\ell};\mathbb{Q})\cong\mathbb{Q}$ since $Y_{\ell}$ is diffeomorphic to $S^{2}\times S^{3}$, and by Theorem 1.1, $H_{3}(W;\mathbb{Q})\cong\mathbb{Q}$. As the Milnor fiber $V_{\ell}$ is homotopy equivalent to the wedge sum of $2\ell-1$ 33-spheres by [Mil68, Theorems 6.5 and 9.1], we obtain $H_{3}(V_{\ell};\mathbb{Q})\cong\mathbb{Q}^{2\ell-1}$. Moreover, this homotopy type implies that $H_{4}(V_{\ell},Y_{\ell})\cong H^{2}(V_{\ell})=0$, which concludes that the induced map $H_{3}(Y_{\ell};\mathbb{Q})\rightarrow H_{3}(V_{\ell};\mathbb{Q})$ by the inclusion is injective; so is the first map in (4.1). To sum up, the sequence (4.1) becomes
(4.2) $\mathbb{Q}\hookrightarrow\mathbb{Q}^{2\ell-1}\oplus H_{3}(W_{\ell,1};\mathbb{Q})\rightarrow\mathbb{Q}\xrightarrow{\varphi}\mathbb{Q}$
Now there are two possibilities of $\varphi$: either it is the zero-map or an isomorphism. In the former case, the sequence (4.2) becomes a split short exact sequence. Hence, we have
$\mathbb{Q}^{2\ell-1}\oplus H_{3}(W_{\ell,1};\mathbb{Q})\cong\mathbb{Q}^{2},$
which contradicts our assumption about $\ell$. Similarly, in the latter case, we have
$\mathbb{Q}^{2\ell-1}\oplus H_{3}(W_{\ell,1};\mathbb{Q})\cong\mathbb{Q},$
again a contradiction. ∎
Remark 4.2.
As mentioned above, $Y_{\ell}$ is diffeomorphic to $S^{2}\times S^{3}$ for every $\ell$. In particular, there are no topological obstructions to the existence of smooth cobordisms between $Y_{\ell}$’s. In this sense, Theorem 4.1 highlights intriguing differences between symplectic and smooth cobordism categories.
Remark 4.3.
The same result as Theorem 4.1 also holds for the $(4d+1)$-dimensional Brieskorn manifolds
$Y_{\ell}^{4d+1}=\{(z_{1},z_{2},\ldots,z_{2d+2})\in\mathbb{C}^{2d+2}\mid z_{1}^{2\ell}+z_{2}^{2}+\cdots+z_{2d+2}^{2}=0\}\cap S^{4d+3}$
for any $d\geq 2$. This follows from essentially the same argument as in the above proof, making use of the middle-dimensional homology of a hypothetical exact symplectic cobordism. Note that $Y_{1}^{4d+1}$, with its canonical contact structure, is contactomorphic to $(ST^{*}S^{2d+1},\xi_{\operatorname{can}})$, just as in the case $d=1$. Moreover, as shown in [Ueb16, Proposition 1], the diffeomorphism types of $Y_{\ell}^{4d+1}$ depend on $\ell$ modulo $4$; in particular, $Y_{\ell}^{4d+1}$ with $\ell\equiv 1\pmod{4}$ is diffeomorphic to $ST^{*}S^{2d+1}$. Thus, in this case there are no topological obstructions to the existence of an exact symplectic cobordism.
Remark 4.4.
Here are some remarks from the perspective of Floer theory.
  1. (1)
    The $5$-dimensional Brieskorn manifolds $Y_{\ell}$ have provided particularly interesting examples of exotic contact structures. In [Ueb16, Corollary 4], it is proved that $Y_{\ell}$’s are pairwise non-contactomorphic using positive (non-equivariant) symplectic homology, while they have the same positive equivariant symplectic homology [KvK16, Appendix A]; see also [KvK16, Question 6.10].
  2. (2)
    As an alternative approach to Theorem 4.1, one might try to show that there are no exact symplectic embeddings from the Milnor fiber $V_{\ell}$ to the Milnor fiber $V_{\ell^{\prime}}$ if $\ell^{\prime}$ is smaller than $\ell$; this implies that there are no cobordisms from $Y_{\ell}$ to $Y_{\ell^{\prime}}$. Unfortunately, most of embedding obstructions in the literature are only available for domains in $\mathbb{R}^{2n}$ or more generally for domains with vanishing symplectic homology, whereas Milnor fibers we are interested in have nonvanshing symplectic homology; see [KvK16, Theorem 6.3]. An obstruction applicable to domains with nonvanishing symplectic homology comes from the notion of dilation recently introduced in [Zho22]. Direct computations using [Zho22, Proposition 5.14] and [KvK16, Section 5.3.1] however show that the Milnor fiber $V_{\ell}$ admits the same order of dilations for each $\ell$, so this obstruction does not work in our situation. As for another potential obstruction, we refer the reader to [GS24].
Remark 4.5.
The Brieskorn manifold $Y^{\prime}_{\ell}$ given by
$Y^{\prime}_{\ell}=\{(z_{1},z_{2},z_{3},z_{4})\in\mathbb{C}^{4}\mid z_{1}^{2\ell-1}+z_{2}^{2}+z_{3}^{2}+z_{4}^{2}=0\}\cap S^{7},$
with $\ell\geq 1$, is known to be diffeomorphic to $S^{5}$ (see e.g. [Bri66]) and has the canonical contact structure $\xi^{\prime}_{\ell}$. One can verify that $(Y^{\prime}_{1},\xi^{\prime}_{1})$ is contactomorphic to the standard sphere $(S^{5},\xi_{\mathrm{st}})$. Indeed, the submanifolds of $S^{7}$ defined by
$Y^{\prime}_{1}(t)=\{(z_{1},z_{2},z_{3},z_{4})\in\mathbb{C}^{4}\mid z_{1}+t(z_{2}^{2}+z_{3}^{2}+z_{4}^{2})=0\}\cap S^{7},$
with $t\in[0,1]$, are all diffeomorphic to $S^{5}$, and $Y^{\prime}_{1}(0)$ and $Y^{\prime}_{1}(1)$ equipped with the contact structures induced by $(S^{7},\xi_{\mathrm{st}})$ agrees with $(S^{5},\xi_{\mathrm{st}})$ and $(Y^{\prime}_{1},\xi^{\prime}_{1})$, respectively; hence, identifying each $Y^{\prime}_{1}(t)$ with $S^{5}$, the Gray stability theorem shows that they are contactomorphic. It is shown in [Ust99] that for $2\ell-1=\pm 1$ (mod $8$), $Y^{\prime}_{\ell}$’s are pairwise non-contactomorphic to each other.
We can show that there exists an exact symplectic cobordism from $(Y^{\prime}_{1},\xi^{\prime}_{1})$ to $(Y^{\prime}_{\ell},\xi^{\prime}_{\ell})$ for each $\ell>1$ but no such a cobordism in the reverse direction, using the same idea as the above proof coupled with the uniqueness result on the diffeomorphism types of symplectically aspherical fillings of $(S^{5},\xi_{\mathrm{st}})$ by Eliashberg–Floer–McDuff [McD91, Theorem 1.5]. Moreover, this also proves the non-existence of exact symplectic cobordisms from $(Y^{\prime}_{\ell},\xi^{\prime}_{\ell})$ to $(Y_{1},\xi_{1})$ or from $(Y_{\ell},\xi_{\ell})$ to $(Y^{\prime}_{1},\xi^{\prime}_{1})$ for $\ell>2$ although there are smooth cobordisms between them.
4.2. Topology of closed symplectic $6$-manifolds
Here we will provide another application of Theorem 1.1 in dimension $6$.
Given a closed symplectic $2n$-manifold $(M,\Omega)$, with $\Omega]\in H^{2}(M;\mathbb{Z})$, having a symplectic hyperplane section $\Sigma$, there is a Liouville vector field $V$, defined near the boundary of a symplectic tubular neighborhood $N(\Sigma)$ of $\Sigma$ in $M$, which points inward along the boundary. We equip this boundary $\partial N(\Sigma)$ with the contact structure defined by $\alpha=(i_{V}\Omega)|_{T\partial N(\Sigma)}$. Then, this contact manifold is contactomorphic to the Boothby–Wang bundle associated with $(\Sigma,\Omega_{\Sigma})$, where $\Omega_{\Sigma}$ denotes the restriction of $\Omega$ to $T\Sigma$. By [DL19, Lemma 2.2], the complement $M\setminus\operatorname{Int}(N(\Sigma))$ with $\Omega|_{M\setminus\operatorname{Int}(N(\Sigma))}$ serves as an exact symplectic filling of the contact manifold $(\partial N(\Sigma),\operatorname{Ker}(\alpha))$.
Theorem 4.6.
Let $(M,\Omega)$ be a closed symplectic $6$-manifold with integral symplectic form. If $(M,\Omega)$ admits a symplectic hyperplane section symplectomorphic to $(\mathbb{CP}^{1}\times\mathbb{CP}^{1},\omega_{\operatorname{FS}}\oplus\omega_{\operatorname{FS}})$, then $M$ is diffeomorphic to the complex $3$-dimensional quadric
$Q^{3}=\{(z_{0}:\ldots:z_{4})\in\mathbb{CP}^{4}\mid z_{0}^{2}+\cdots+z_{4}^{2}=0\}.$
The proof of the theorem relies on a classification theorem of certain $6$-manifolds.
Theorem 4.7 (Wall [Wal66] and Jupp [Jup73]).
The following topological invariants determine the diffeomorphism types of smooth, closed, oriented, simply connected $6$-manifolds $M$ with torsion-free homology:
  • the two cohomology groups $H^{2}(M;\mathbb{Z})$ and $H^{3}(M;\mathbb{Z})$;
  • the triple cup product $\mu\colon H^{2}(M;\mathbb{Z})\otimes H^{2}(M;\mathbb{Z})\otimes H^{2}(M;\mathbb{Z})\rightarrow\mathbb{Z}$, $\mu(a,b,c)=(a\smile b\smile c)[M]$;
  • the second Stiefel–Whitney class $w_{2}(M)\in H^{2}(M;\mathbb{Z}_{2})\cong H^{2}(M;\mathbb{Z})\otimes\mathbb{Z}_{2}$;
  • the first Pontryagin class $p_{1}(M)\in H^{4}(M;\mathbb{Z})\cong\operatorname{Hom}(H^{2}(M;\mathbb{Z}),\mathbb{Z})$.
Proof of Theorem 4.6.
In what follows, we show the invariants given in Theorem 4.7 are completely determined by the information of the normal bundle of $\Sigma$ in $M$, which is isomorphic to the one of $Q^{2}\coloneqq\mathbb{CP}^{1}\times\mathbb{CP}^{1}$ in $Q^{3}$. Notice that the symplectic manifold $(M,\Omega)$ can be the quadric $Q^{3}$ with the symplectic form induced by the Fubini–Study form on $\mathbb{CP}^{4}$ since $(\mathbb{CP}^{1}\times\mathbb{CP}^{1},\omega_{\operatorname{FS}}\oplus\omega_{\operatorname{FS}})$ appears as a hyperplane section in the sense of complex geometry.
Recall that the Boothby–Wang bundle associated with $(\mathbb{CP}^{1}\times\mathbb{CP}^{1},\,\omega_{\operatorname{FS}}\oplus\omega_{\operatorname{FS}})$ is contactomorphic to $(ST^{*}S^{3},\xi_{\operatorname{can}})$. Hence, removing a symplectic tubular neighborhood $N(\Sigma)$ of $\Sigma$, we obtain an exact symplectic filling $W$ of $(ST^{*}S^{3},\xi_{\operatorname{can}})$, which is diffeomorphic to $DT^{*}S^{3}$ by Theorem 1.1. Since $M=W\cup N(Q^{2})$ and the spaces $W$, $N(Q^{2})$, and $W\cap N(Q^{2})=ST^{*}S^{3}$ are simply connected, it follows that $M$ is simply connected as well. To determine the homology and cohomology of $M$, note that $H_{k}(M;\mathbb{Z})$ and $H^{k}(M;\mathbb{Z})$ are straightforward to compute for all $k\neq 3$ using the exact sequences of the pairs $(M,W)$ and $(M,N(Q^{2}))$. For the case $k=3$, consider the homology exact sequence of $(M,N(Q^{2}))$:
$H_{3}(N(Q^{2});\mathbb{Z})\to H_{3}(M;\mathbb{Z})\to H_{3}(M,N(Q^{2});\mathbb{Z})\to H_{2}(N(Q^{2});\mathbb{Z})\to H_{2}(M;\mathbb{Z}),$
which is equivalently written as
$0\;\to\;H_{3}(M;\mathbb{Z})\;\to\;\mathbb{Z}\;\to\;\mathbb{Z}\oplus\mathbb{Z}\;\to\;\mathbb{Z}.$
The group $H_{2}(N(Q^{2});\mathbb{Z})$ is generated by the two ruling curves of $Q^{2}\cong\mathbb{CP}^{1}\times\mathbb{CP}^{1}$, while $H_{2}(M;\mathbb{Z})$ is generated by one such curve. Thus the map $H_{2}(N(Q^{2});\mathbb{Z})\to H_{2}(M;\mathbb{Z})$ has kernel isomorphic to $\mathbb{Z}$, which is free, and hence $H_{3}(M,N(Q^{2});\mathbb{Z})\to H_{2}(N(Q^{2});\mathbb{Z})$ is injective. Therefore $H_{3}(M;\mathbb{Z})\cong H^{3}(M;\mathbb{Z})=0$. We conclude that
$H_{k}(M;\mathbb{Z})\cong H_{k}(Q^{3};\mathbb{Z})\quad\text{and}\quad H^{k}(M;\mathbb{Z})\cong H^{k}(Q^{3};\mathbb{Z})\quad\text{for all }k.$
In order to see the rest of the invariants in Theorem 4.7, we observe generators of $H_{2}(M;\mathbb{Z})$, $H_{4}(M;\mathbb{Z})$ and $H^{2}(M;\mathbb{Z})$. In view of the above argument, the image of $\mathbb{CP}^{1}]\in H_{2}(\mathbb{CP}^{1};\mathbb{Z})$ under the map induced by the inclusion $j\colon\mathbb{CP}^{1}\hookrightarrow\mathbb{CP}^{1}\times\{\mathrm{pt}\}\xhookrightarrow{i}M$ gives rise to a generator of $H_{2}(M;\mathbb{Z})\cong\mathbb{Z}$. As for the cohomology group $H^{2}(M;\mathbb{Z})$, since
$\Omega](j_{*}[\mathbb{CP}^{1}])=[\omega_{\operatorname{FS}}]([\mathbb{CP}^{1}])=1,$
the cohomology class $\Omega]$ is a primitive element and hence a generator of the group. Similarly, the intersection
$\Sigma]\cdot j_{*}[\mathbb{CP}^{1}]=1$
shows that $\Sigma]=i_{*}[\mathbb{CP}^{1}\times\mathbb{CP}^{1}]$ generates $H_{4}(M;\mathbb{Z})$. Notice that these generators are determined by the information of the neighborhood $N(\Sigma)$ of $\Sigma$.
Now the triple tensor product $\Omega]^{\otimes 3}$ generates $H^{2}(M;\mathbb{Z})^{\otimes 3}$, and the fact that $\operatorname{PD}([\Omega])=[\Sigma]$ gives
$\mu([\Sigma]^{\otimes 3})$ $=([\Omega]\smile[\Omega]\smile[\Omega])([M])=([\Omega]\smile[\Omega])([\Sigma])$
$=([\omega_{\operatorname{FS}}\oplus\omega_{\operatorname{FS}}]\smile[\omega_{\operatorname{FS}}\oplus\omega_{\operatorname{FS}}])([\mathbb{CP}^{1}\times\mathbb{CP}^{1}]).$
As complex vector bundles, $i^{*}TM$ and $j^{*}TM$ split into
$i^{*}TM\cong T(\mathbb{CP}^{1}\times\mathbb{CP}^{1})\oplus\mathcal{O}_{\mathbb{CP}^{1}\times\mathbb{CP}^{1}}(1)\quad{\mathrm{and}}\quad j^{*}TM\cong\mathcal{O}_{\mathbb{CP}^{1}}\oplus\mathcal{O}_{\mathbb{CP}^{1}}(1)\oplus\mathcal{O}_{\mathbb{CP}^{1}}(2),$
respectively, which is evidently independent of $M\setminus N(\Sigma)$. According to the naturality of characteristic classes, $w_{2}(M)$ and $p_{1}(M)$ are completely determined by those pull-back bundles, which completes the proof. ∎
Remark 4.8.
To the best of the authors’ knowledge, it is a difficult problem to determine whether a closed symplectic $6$-manifold $(M,\Omega)$ satisfying the assumption of Theorem 4.6 is symplectomorphic to the $3$-dimensional quadric $Q^{3}$ equipped with the standard Kähler form. Nevertheless, a result of Li and Ruan [LR09] reveals a symplectic feature of $(M,\Omega)$: by [LR09, Corollary 4.6], $(M,\Omega)$ is symplectically uniruled, as defined in terms of Gromov–Witten invariants.
Remark 4.9.
A similar result to Theorem 4.6 holds for $(\mathbb{CP}^{2},\omega_{\operatorname{FS}})$: if a closed symplectic symplectic $6$-manifold $(M,\Omega)$ with $\Omega]\in H^{2}(M;\mathbb{Z})$ contains $(\mathbb{CP}^{2},\omega_{\operatorname{FS}})$ as a symplectic hyperplane section, then $M$ is diffeomorphic to $\mathbb{CP}^{3}$. To see this, by using the result of Eliashberg–Floer–McDuff [McD91] mentioned in Remark 4.5, one can check that such a manifold $M$ is simply connected and has the homology of $\mathbb{CP}^{3}$. Then, the rest of the proof goes along the same line as the one for Theorem 4.6.
Acknowledgements
The authors thank Otto van Koert and Kai Zehmisch for valuable comments, and in particular Zhengyi Zhou for pointing out a gap in the first version of this paper. The first author was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2021R1F1A1060118, RS-2025-23524132). The second author was supported by Japan Society of Promotion of Science KAKENHI Grant numbers JP20K22306, JP22K13913, JP24H00182.